OpenStudy (anonymous):

6+y=7 6+y-6 left 7-6 right now how do I write the left and the right side of the relation. I know I add 7 and the answer is x=1

OpenStudy (anonymous):

I meant y=1

OpenStudy (anonymous):

-6 +6+y = 7-6 that will equal y=7-6 . How do I write y=1? Do I need to add ?

OpenStudy (anonymous):

Simplify by y=7-6 and y=1?

OpenStudy (anonymous):

yes... that is perfectly correct :)